SX AO Science

04:06
  • Configuring for AO
  • Seeing is somewhat variable, and at about 1.2, but this is bin2, so, we should be fine.
04:11
  • Restarted the IDL and fastlink processes, and restarted the arbitrator, to fix the mirror GUI.
04:17
  • Pointing
04:21
  • Seeing 1.06"
  • Collimating
  • Working well
  • Going with previous gains, 2.00 0.90 0.40
  • Measured brightness 13.2
  • I think that I heard them say that there was a MOS error.
  • 043023
04:32
  • It was a MOS error
  • Switching to LUCI2, until Dave can get home and fix it.
06:02
  • MOS is KAPUT! No LUCI1 Tonight.

-- %USERSIG{GregoryTaylor - 2018-02-04}%
Topic revision: r2 - 04 Feb 2018, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback